Suppose price of commodity X is Rs. 1,000 and price of commodity is Rs. 500 and a consumers has Rs. 20,000 to spent per month on good X and Y. a Sketch the budget constraint 5. Assume that he splits his income equally between X & Y. Show where the consumer ends up on the budget constraints. c Suppose that income rises from Rs. 20,000 to Rs. 40,000 sketch the new budget constraint. d. Assume that he again splits total budget equally on two goods Show where the consumer ends up on the new budget constraint. 


1
Expert's answer
2021-03-21T20:04:36-0400
"1000x+500y=20000"

a.



b.

"1000x=10000"


"x=10"


"500y=10000"

"y=20"

Checking the budget constraint


"1000\times10+500\times20=20000"


c.


"1000x+500y=40000"


d.